# von Neumann entropy and quantum version of thermodynamic entropy

### v2

@@ Abstract
The debate whether the von Neumann (VN) entropy is a suitable quantum version of the thermodynamic (TH) entropy has a long history. In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be addressed convincingly. In a broader context, we here analyze in detail whether and when the VN entropy is the same or equivalent to the TH entropy. We find that the VN entropy of a large isolated system is equivalent to its TH entropy if the system is in a mixed state described by a microcanonical (MC) ensemble. This equivalence also holds for open system and subsystem if the full system (i.e. system-bath or subsystem-bath combination) is in a mixed state described by an MC ensemble or in a typical pure state (or if the Eigenstate Thermalization Hypothesis, ETH, holds for the full system). Unlike isolated and open system, demonstrating the equivalence for a subsystem is not straightforward. We take a one-dimensional spin-1/2 model to numerically show that, if the full system is in a pure state, then the VN entropy of a subsystem is equivalent to its TH entropy if the ETH holds for the full system. We also demonstrate that this equivalence always works (irrespective of whether the ETH holds) if the full system is in a mixed state describe by an MC ensemble.
preprint: APS/123-QED
I Introduction
@@
Since the development of quantum mechanics, there has been an enormous progress in our understanding of how the quantum many-body systems behave. This progress saw some foundational questions being asked and at-least partially answered. One of such questions is on whether and how the statistical mechanics works for the quantum many-body systems[1, 2, 3]. This gives rise to the important field of quantum statistical mechanics (QSM). The thermodynamic (TH) entropy is a very important state function which helps us understand how a system approaches equilibrium besides playing a crucial role in studying equilibrium properties of a system [4]. The usefulness of the TH entropy naturally calls for its appropriate quantum counterpart in QSM.
The TH entropy, as defined by Clausius in the context of second law of thermodynamics, is a state function which takes the maximum value for a system in the equilibrium under fixed macroscopic conditions [4, 5]. For an isolated system, this entropy never decreases when the system undergoes some process. Thus any entropy defined in the similar spirit for a quantum system must be a state function and should increase (or remain constant) when the system undergoes a process.
Boltzmann, and later Gibbs, showed us how the TH entropy can be calculated from the microscopic details using statistical approaches [4, 6]. For an isolated system, Boltzmann gave a formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates). For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms of an appropriate probability distribution function in phase space. Although the quantum versions of these statistical entropies for the isolated and open systems in equilibrium are somewhat intuitively clear [7, 5], one needs convincing justifications, which we discuss in Sec. II. The question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system. In this work, we will discuss this issue in detail.
As noted, for a quantum system in thermal equilibrium, writing down the quantum version of the Gibbs (GB) entropy is somewhat easy (see, Secs. II.3 and III.1). But when a quantum system is in a generic mixed state and not in equilibrium, what would be the value of entropy? From thermodynamical considerations, von Neumann (VN) gave a formula for the entropy of a quantum ensemble (representing a given mixed state) [8]. This entropy, known as VN entropy, attracted a great deal of attention and a lot of criticism in regards to its connection to the TH entropy. We show in this work that, for isolated and open quantum system or for a subsystem of a (much) larger quantum system, the VN entropy is equivalent to the TH entropy if the quantum system (or subsystem) is large and in thermal equilibrium. We largely exclude from our discussions the quantum systems that do not thermalize or systems that are not in equilibrium. We here, though, deliberate on two main criticisms on the VN entropy, related to the time-invariance and subadditivity, and argue that these criticisms can be avoided or addressed convincingly.
We outline here the logical steps we follow in this work while establishing the said equivalence between the VN and TH entropies and finding the conditions under which the equivalence works. First, we identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: the dimension of relevant subspace in place of the volume of accessible phase-space, and in Gibbs’ case: the energy eigenvalues in place of the continuous classical energy). In the last step of establishing the equivalence, we check if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e. the reinterpreted statistical entropy in terms of quantum variables). To better understand the equivalence, one should also check if the VN entropy inherits the important properties of the equilibrium TH entropy. Accordingly, wherever appropriate, we briefly deliberate if the VN entropy is a state function, has the additivity property and whether it increases for isolated system due to an irreversible process.
We organize our paper in the following way. In the next section (Sec. II), we review the thermodynamic and statistical entropies, and discuss the condition of equivalence between these entropies. This is followed by a discussion on the suitable quantum versions of the statistical entropies. In Sec. III, we discuss von Neumann entropy and some reservations about the entropy when it comes to representing the thermodynamic entropy. Next, in Sec. IV, we check for three different cases (isolated system, open system and subsystem) if the von Neumann entropy is equivalent to the thermodynamic entropy. We conclude our work in Sec. V.
II Thermodynamic and statistical entropies
@@
In this section we first briefly review the contexts and definitions of the thermodynamic (TH) entropy and the statistical (ST) entropies (both Boltzmann’s and Gibbs’ versions). Next we try to understand under what conditions these entropies (TH and ST) are equivalent. At the end this section, we discuss the reinterpretations of the statistical entropies in terms of the quantum variables.
II.1 Entropy: from Clausius to Boltzmann and Gibbs
@@
The Boltzmann’s version of statistical entropy is defined by the number of accessible microstates $\Omega$ of a system (with a fixed energy). For a gaseous system, this number is proportional to the phase-space volume available to the system in the given macrostate. Here the statistical entropy is defined as $S_{BN}=k_{B}\ln{\Omega}$ ) with $k_{B}$ being the Boltzmann (BN) constant [4, 9, 7]. By dividing the accessible phase-space as per the values of the well-defined macrovariables, one can in-fact calculate the BN entropy for each such division [7, 5]. This way one can even make the BN entropy useful in studying a system approaching equilibrium.
@@
II.2 Conditions of equivalence between thermodynamic and statistical entropies
To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence. A common way of showing the equivalence between the TH and BN entropies of a system with a given energy is to bring the system with energy $E_{1}$ in contact with another system with energy $E_{2}$ [6]. Before the contact, the systems were in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat. Although the systems exchange heat, the total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems. In the physical setup to establish the equivalence, one generally employs the principle of equal a priori probabilities. According to this premise, a thermodynamic system in equilibrium with a given energy can be found in any of the available microstates with equal probability. Here the number of available microstates is a function of the system’s energy. If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then in this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$. We here assume that the two systems can independently be in any of their available microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored). The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values (i.e., over all possible energy distributions). As per the premise, the composite system in equilibrium is equally likely to be in any of the $M(E)$ number of microstates. When the systems are thermodynamically large, $M(E)$ will be almost equal to $\Omega(\overline{E}_{1})$, i.e., $M(E)\approx\Omega(\overline{E}_{1})$, where $E^{\prime}_{1}=\overline{E}_{1}$ maximizes the function $\Omega(E^{\prime}_{1})$. This means that the composite system, after settling in a new equilibrium state, will spend most of its time in the microstates corresponding to the energy distribution $\overline{E}_{1}$ and $E-\overline{E}_{1}$ for the first and second systems, respectively.
@@ II.2 Conditions of equivalence between thermodynamic and statistical entropies
II.3 Quantum version of statistical entropy
@@
Once we have agreement on the quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies. To get the quantum version of the BN (qBN) entropy, we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian and {|$\{\ket{n}\}$ be the corresponding set of eigenkets. As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$. Subsequently, guided by our understanding of the BN entropy, we can write its quantum version ($S_{qBN}$) in the following way:
$S_{BN}\to S_{qBN}=k_{B}\ln D,$ (1)
@@
where $D$ is the number of eigenvalues within the shell. It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes (the largest dimension) [7, 5]. For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
@@
To get the quantum version of the Gibbs (qGB) entropy, in place of continuous classical energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function. Here we assume that the system’s interaction with the reservoir is weak (or it had interaction with the reservoir in the past), so that we still can effectively describe the dynamics of the system by its Hamiltonian. As before, let $\{E_{n}\}$ be the set of eigenvalues of the Hamiltonian of the (open) system. In the spirit of the Gibbs distribution for the classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. Subsequently, we can write the quantum version of the GB entropy in the following way:
$S_{GB}\to S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}.$ ) . (2)
We here briefly note that the probabilities $p_{n}$ are the eigenvalues of the density matrix $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ is the Hamiltonian of the (open) system. Using the density matrix $\rho_{\beta}$, the quantum Gibbs (qGB) entropy in Eq.2 can also be expressed in a basis-independent form, as discussed in Sec.III.1. It may be emphasized that the relevant density matrix is found here indirectly: starting from a classical (open) system, we first find the probability distribution and then reinterpreting the quantities in terms of the quantum variables, we get the density matrix. Instead, it is also possible to get the density matrix $\rho_{\beta}$ directly, starting with a quantum system [10, 11]. In this work, we prefer the first method since it directly reveals when and how the TH entropy is equivalent to the qGB entropy.
III von Neumann Entropy
@@
$S_{VN}=-k_{B}\text{Tr}(\rho\ln\rho).$ (3)
This formula attracted both the applause and criticism alike [12, 13, 14, 15, 16]. In the following subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies. Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or can be addressed convincingly.
III.1 As basis-independent representation of quantum statistical entropy
One way to understand the VN entropy is to view it as the basis-independent representation of the quantum statistical entropy of a thermodynamic system in equilibrium. The VN entropy for an appropriate density matrix can give us the qBN or qGB entropy. In fact, while deriving the expression of the thermodynamic (TH) entropy of a subsystem, Landau also came up with the same expression that appears in Eq. 3 as the basis-independent representation of the quantum statistical entropy [9].
For an isolated system, described by a microcanonical (mc) ensemble, we get the following from Eq. 1: $S_{qBN}=-k_{B}\ln D=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=S_{VN}$, where $\rho_{mc}=\frac{1}{D}I$ with $D$ being the dimension of the relevant subspace of the Hilbert space of the system and $I$ is the identity matrix. So we find here that the VN entropy can be viewed as a basis-independent representation of the qBN entropy of an isolated system in thermal equilibrium.
Similarly, for an open system, we get the following from the Eq. 2, $S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=S_{VN}$, where $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$. Here $H$ and $\beta$ are, respectively, the Hamiltonian and the inverse temperature of the open system. We note again that the VN entropy of an open system in equilibrium can be viewed as a basis-independent representation of the qGB entropy.
III.2 As entropy of thermodynamic ensemble

### v3

# Statistical entropy of quantum systems

@@ Abstract
Let $D_{1}$ and $D_{2}$ be the Hilbert space dimensions of two subsystems of a quantum system of total Hilbert space dimension $D=D_{1}D_{2}$. In the thermodynamic limit (with $1\ll D_{1}\ll D_{2}$), we know from the works of Page and Sen that the average von Neumann (VN) entropy of the first subsystem is $\mathbb{E}({S}^{sb}_{VN})=\ln(D_{1})+O(D_{1}/D_{2})$ if the full system is in a random pure state. Here, it is argued that this result can be strengthened for a thermalized quantum system. Consider the subspace $\mathcal{H}_{E}$ of the total Hilbert space corresponding to a narrow shell around the energy $E$. We find that the result of Page and Sen holds for each of these subspaces, that is, the VN entropy, when averaged over the states in $\mathcal{H}_{E}$, is given by $\overline{S}^{sb}_{VN}\approx\ln\widetilde{d}_{1}$, where $\widetilde{d}_{1}$ represents the dimension of the effective Hilbert space of the first subsystem relevant to $\mathcal{H}_{E}$. If $d_{E}=\dim{(\mathcal{H}_{E})}$, we estimate that $\widetilde{d}_{1}=D_{1}^{\gamma}$, where $\gamma=\ln(d_{E})/\ln(D)$. This finding has significant implications, as it suggests an equivalence between the VN entropy and the thermodynamic (TH) entropy of a subsystem within a much larger thermalized quantum system. For completeness, we also discuss in this work the issue of equivalence between the VN entropy and TH entropy for isolated system (as a whole) and open system. For numerical demonstration of our important results, we here consider a one-dimensional spin-1/2 chain with next-nearest neighbor interactions.
Keywords: Entropy of quantum systems, von Neumann entropy, Thermodynamic entropy
1 Introduction
@@
Since the development of quantum mechanics, there has been an enormous progress in our understanding of how the quantum many-body systems behave. This progress saw some foundational questions being asked and at-least partially answered. One of such questions is on whether and how the statistical mechanics works for the quantum many-body systems[10, 4, 27]. This gives rise to the important field of quantum statistical mechanics (QSM). The thermodynamic (TH) entropy is a very important state function which helps us understand how a system approaches equilibrium besides playing a crucial role in studying equilibrium properties of a system [16]. The Boltzmann (BN) entropy and Gibbs (GB) entropy are two well-established statistical measures that represent the thermodynamic (TH) entropy of classical systems. The success of these statistical entropies inspires the search for similar statistical quantities that can accurately represent the TH entropy of quantum systems. In this regard, a key question is whether the von Neumann (VN) entropy is a suitable candidate for this role. Against this backdrop, here we explore some key challenges and open questions in the field of QSM. Our primary contributions are as follows.
We argue here that the landmark result of Page [22] and Sen [30] can be strengthened for a thermalized system. Their result roughly shows that the von Neumann (VN) entropy of a subsystem, when the full system is in a random pure state, is equal to the logarithm of the dimension of the subsystem’s Hilbert space. For a thermalized system, we show that a result equivalent to that of Page and Sen holds for any subspace of the full Hilbert space associated with a narrow shell around some energy $E$.
We discuss the above result in a broader context. As mentioned in the beginning, it is of fundamental interest to know whether the VN entropy is a suitable statistical measure of the TH entropy of a quantum system. The above-mentioned result answers an important part of this general question: it is argued that this result implies an equivalence between the VN entropy and the TH entropy for a subsystem when the full (thermalized) system is in a (random) pure state associated with a narrow energy shell. This question of equivalence is also addressed using the notion of canonical typicality [26, 11].
We also address the question of equivalence between the VN entropy and the TH entropy of a subsystem when the full system is in a mixed state. We emphasize that understanding the equivalence between the VN entropy and the TH entropy of a subsystem is crucial in the context of quantum thermalization, as thermalization is essentially defined with respect to a subsystem (or a local operator) [19, 5, 33]. For completeness and better understanding, the same question of equivalence is also addressed for open system (with very weak interaction with the environment) and isolated system (as a whole). For demonstration of the nontrivial results, we also perform numerical analysis on a spin-1/2 chain with next-nearest neighbor interactions.
Since understanding the equivalence between the VN entropy and the TH entropy (in any of the cases mentioned above) requires evaluating both the entropies separately, we first, in Sec. 2, deliberate on how one can quantify the TH entropy of a quantum system. We also provide a detailed discussion on the conditions under which this quantification holds. Next, in Sec. 3, we have a general discussion on the VN entropy and address two major objections, related to time invariance and subadditivity properties, when compared to the TH entropy in a quantum system. In Sec. 4, we discuss the issue of equivalence between the VN entropy and the TH entropy separately for three cases: isolated system, open system and subsystem of a larger system. We conclude our work in Sec. 5.
2 Thermodynamic and statistical entropies
@@
In this section we first briefly review the thermodynamic (TH) entropy and the statistical (ST) entropies (both Boltzmann’s and Gibbs’ versions). Next we try to understand under what conditions these entropies (TH and ST) are equivalent. At the end of this section, we discuss the reinterpretations of these statistical entropies in terms of the quantum variables.
2.1 Entropy: from Clausius to Boltzmann and Gibbs
@@
The Boltzmann’s version of statistical entropy is defined by the number of accessible microstates $\Omega$ of a system (with a fixed energy). For a gaseous system, this number is proportional to the phase-space volume available to the system in the given macrostate. Here the statistical entropy is defined as $S_{BN}=\ln{\Omega}$ (in this paper we set the Boltzmann constant $k_{B}=1$) [16, 17, 12]. By dividing the accessible phase-space as per the values of the well-defined macrovariables, one can in-fact calculate the BN entropy for each such division [12, 19]. This way one can even make the BN entropy useful in studying a system approaching equilibrium.
@@
2.2 Conditions of equivalence between thermodynamic and statistical entropies
To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence. A common way of showing the equivalence between the TH and BN entropies of a system with a given energy is to bring the system with energy $E_{1}$ in contact with another system with energy $E_{2}$ [23]. Before the contact, the systems were individually in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat. Although the systems exchange heat, the total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems. In the physical setup to establish the equivalence, one generally employs the principle of equal a priori probabilities. According to this premise, a thermodynamic system in equilibrium with a given energy can be found in any of the available microstates with equal probability. Here the number of available microstates is a function of the system’s energy. If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then in this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$. We here assume that the two systems can independently be in any of their available microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored). The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values (i.e., over all possible energy distributions). As per the premise, the composite system in equilibrium is equally likely to be in any of the $M(E)$ number of microstates. When the systems are thermodynamically large, $M(E)$ will be almost equal to $\Omega(\overline{E}_{1})$, i.e., $M(E)\approx\Omega(\overline{E}_{1})$, where $E^{\prime}_{1}=\overline{E}_{1}$ maximizes the function $\Omega(E^{\prime}_{1})$. This means that the composite system, after settling in a new equilibrium state, will spend most of its time in the microstates corresponding to the energy distribution $\overline{E}_{1}$ and $E-\overline{E}_{1}$ for the first and second systems, respectively.
@@ 2.2 Conditions of equivalence between thermodynamic and statistical entropies
2.3 Reinterpreting statistical entropies for quantum systems
@@
Once we have agreement on the quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies. To get the quantum version of the BN entropy (qBN entropy), we prepare a system in such a way that its energy lies in the narrow energy shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [19]. Let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian and {|$\{\ket{n}\}$ be the corresponding set of eigenkets. As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$. Subsequently, guided by our understanding of the BN entropy, we can write its quantum version ($S_{qBN}$) in the following way:
$S_{BN}\to S_{qBN}=\ln d_{E},$ (1)
@@
where $d_{E}$ is the number of eigenkets (microstates) associated with the shell. It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes [12, 19]. For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to $d_{E}$.
@@
To get the quantum version of the Gibbs (qGB) entropy, in place of continuous classical energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function. Here we assume that the system’s interaction with the reservoir is weak (or it had interaction with the reservoir in the past), so that we still can effectively describe the dynamics of the system by its Hamiltonian. As before, let $\{E_{n}\}$ be the set of eigenvalues of the Hamiltonian of the (open) system. In the spirit of the Gibbs distribution for the classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. Subsequently, we can write the quantum version of the GB entropy (qGB entropy) in the following way:
$S_{GB}\to S_{qGB}=-\sum_{n}p_{n}\ln{p_{n}}.$ (2)
We here briefly note that the probabilities $p_{n}$ are the eigenvalues of the density matrix $\rho_{\beta}=\frac{e^{-\beta H}}{\mathrm{Tr}(e^{-\beta H})}$, where $H$ is the Hamiltonian of the (open) system. Using the density matrix $\rho_{\beta}$, the quantum Gibbs (qGB) entropy in Eq.2 can also be expressed in a basis-independent form, which looks the same as the VN entropy formula as appears in Eq. 3. It may be emphasized that the quantum statistical entropy (qBN or qGB) of an isolated or open system can be put into a basis-independent form, appearing exactly the same as the VN entropy formula [17]. But to show the equivalence between the VN entropy and the TH entropy (represented by qBN or qGB entropy), one should evaluate them independently. In Sec. 4, we calculate them separately across all three cases mentioned earlier.
3 von Neumann Entropy
@@
$S_{VN}=-\mathrm{Tr}(\rho\ln\rho).$ (3)
This formula attracted both the applause and criticism alike [31, 13, 14, 7, 35]. In the following subsections, we first discuss two different perspectives to understand how one can arrive at this formula and what it actually quantifies. Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or can be addressed convincingly.
3.1 Entropy of thermodynamic ensemble